OpenStudy (anonymous):

7x -4y=20 for y when x=5

OpenStudy (ranga):

substitute x = 5 in the equation and solve for y.

OpenStudy (anonymous):



OpenStudy (ranga):

Why do you multiply 4 and 20?

OpenStudy (ranga):

7x -4y=20 for y when x=5 Put x = 5 in the equation: 7*5 - 4*y = 20 35 - 4*y = 20 subtract 20 from both sides: 35 - 20 - 4*y = 0 15 - 4*y = 0 add 4*y to both sides 15 = 4*y divide both sides by 4 15/4 = y y = 15/4 or 3&3/4 or 3.75

OpenStudy (anonymous):

cool thanks